OpenStudy (keyonsay):

A punch recipe requires 4/5 of a cup of pineapple juice for every 2 1/2 cups of soda. What is the unit rate of soda to pineapple juice in the punch?

OpenStudy (anonymous):

\[2\tfrac{1}{2}:\frac{4}{5}\] but i guess it wants you do divide \[\frac{5}{2}\div\frac{4}{5}\]

OpenStudy (keyonsay):

THANKS :DDD

OpenStudy (anonymous):

yw